# On the $\Gamma$-limit of weighted fractional energies

### v1

@@ 2.2. Compactness and characterization
$\|\tau_{h}\tilde{u}^{n}-\tilde{u}^{n}\|_{L^{p}(\mathbb{R}^{d})}\leq C(d,p,M,\Omega,f)\|h\|^{s_{n}}$ (3.14)
for all $h\in\mathbb{R}^{d}$ such that $\|h\|\leq 1$. Thus, owing to (3.1) and to (3.14), we can apply theorem 2.1 and find $u\in L^{p}(\mathbb{R}^{d})$ such that, up to a subsequence, $\tilde{u}^{n}\to u$ in $L^{p}(\mathbb{R}^{d})$. Furthermore, since $\tilde{u}^{n}=0$ for all $n\in\mathbb{N}$ on $\mathbb{R}^{d}\setminus\Omega$, we also have that $u=0$ on $\mathbb{R}^{d}\setminus\Omega$. Finally, letting $n\to\infty$ in (3.14), we have
$\|\tau_{h}u-u\|_{L^{p}(\mathbb{R}^{d})}\leq C(d,p,M,\Omega,f)\|h\|$
for all $h$ with $\|h\|\leq 1$, so that $u\in W^{1,p}(\mathbb{R}^{d})$ for $p>1$ or $u\in BV(\mathbb{R}^{d})$ for $p=1$ by theorem 2.2. Since $u=0$ on $\mathbb{R}^{d}\setminus\Omega$ and $\Omega$ has Lipschitz boundary, we get that $u|_{\Omega}\in W^{1,p}_{0}(\Omega)$ for $p>1$, or $u|_{\Omega}\in BV_{0}(\Omega)$ for $p=1$, concluding the proof. ∎
3.2. Proof of theorem 3.1(ii)
@@
is finite. In addition, given $g\colon\mathbb{R}^{d}\to\mathbb{R}$, we let $g_{i}^{a}=\inf_{Q_{i}^{a}}g$. Notice that, whenever $g\in\mathrm{Lip}(\mathbb{R}^{d};(0,\infty))$, then $g_{i}^{a}>0$. Finally, we let $\eta\in C_{c}^{\infty}(B_{1})$ be such that $\eta\geq 0$ and $\int_{B_{1}}\eta\,\mathrm{d}x=1$ and, for every $\varepsilon>0$, we set $\eta_{\varepsilon}(\,\cdot\,)=\varepsilon^{-d}\eta(\,\cdot\,/\varepsilon)$. Accordingly, we let $u_{\varepsilon}=u\ast\eta_{\varepsilon}$ for every $\varepsilon>0$ and $u\in L^{1}_{\rm loc}(\mathbb{R}^{d})$.
@@
Let $p\in[1,\infty)$, $g\in\mathrm{Lip}_{b}(\mathbb{R}^{d};(0,\infty))$. There exist $\varepsilon,\beta,\gamma>0$ with $\varepsilon\ll\beta\ll\gamma$ such that
$\begin{split}\int_{Q_{i}^{(1-\beta)\gamma}}\int_{Q_{i}^{(1-\beta)\gamma}}&\frac{|\tilde{u}_{\varepsilon}(x)-\tilde{u}_{\varepsilon}(y)|^{p}}{\|x-y\|^{d+sp}}(g_{i}^{\gamma})^{2}\,\mathrm{d}x\,\mathrm{d}y\\ &\quad\qquad\leq\int_{Q_{i}^{\gamma}}\int_{Q_{i}^{\gamma}}\frac{|\tilde{u}(x)-\tilde{u}(y)|^{p}}{\|x-y\|^{d+sp}}g(x)g(y)\,\mathrm{d}x\,\mathrm{d}y\end{split}$ (3.15)
@@
for all $z\in B_{\varepsilon}$ and all $i\in\gamma\mathbb{Z}^{d}$. Using the definition of convolution, Jensen’s inequality, changing order of integration, performing the change of variables $x-z=\xi$ and $y-z=\zeta$, owing to (3.16), changing once again order of integration, using the fact that $\eta_{\varepsilon}$ has unit $L^{1}$ norm, and finally owing to the definition of $g_{i}^{\gamma}=\inf_{Q_{i}^{\gamma}}g$, we obtain
$\int_{Q_{i}^{(1-\beta)\gamma}}$ $\int_{Q_{i}^{(1-\beta)\gamma}}\frac{|\tilde{u}_{\varepsilon}(x)-\tilde{u}_{\varepsilon}(y)|^{p}}{\|x-y\|^{d+sp}}(g_{i}^{\gamma})^{2}\,\mathrm{d}x\,\mathrm{d}y$
$\leq\int_{Q_{i}^{(1-\beta)\gamma}}\int_{Q_{i}^{(1-\beta)\gamma}}\int_{B_{\varepsilon}}\frac{|\tilde{u}(x-z)-\tilde{u}(y-z)|^{p}}{\|x-z-(y-z)\|^{d+sp}}\eta_{\varepsilon}(z)(g_{i}^{\gamma})^{2}\,\mathrm{d}z\,\mathrm{d}x\,\mathrm{d}y$
$\leq\int_{B_{\varepsilon}}\int_{Q_{i}^{(1-\beta)\gamma}+z}\int_{Q_{i}^{(1-\beta)\gamma}+z}\frac{|\tilde{u}(\xi)-\tilde{u}(\zeta)|^{p}}{\|\xi-\zeta\|^{d+sp}}\eta_{\varepsilon}(z)(g_{i}^{\gamma})^{2}\,\mathrm{d}\xi\,\mathrm{d}\zeta\,\mathrm{d}z$
$\leq\int_{Q_{i}^{\gamma}}\int_{Q_{i}^{\gamma}}\frac{|\tilde{u}(\xi)-\tilde{u}(\zeta)|^{p}}{\|\xi-\zeta\|^{d+sp}}\,g(\xi)\,g(\zeta)\,\mathrm{d}\xi\,\mathrm{d}\zeta,$
@@
[…]
Proof of theorem 3.1(ii).
@@
otherwise inequality (3.2) is trivially true. Therefore, we can assume the validity of (3.1) which, in turn, implies that $u\in W^{1,p}_{0}(\Omega)$ for $p>1$, or $u\in BV_{0}(\Omega)$ for $p=1$. By lemma 3.5, there exist $\varepsilon,\beta,\gamma>0$ with $\varepsilon\ll\beta\ll\gamma$ such that
@@
$|\displaystyle|\nabla(\tilde{u}^{n})_{\varepsilon}(x)\cdot(x-y)|$$\leq|(\tilde{u}^{n})_{\varepsilon}(x)-(\tilde{u}^{n})_{\varepsilon}(y)|+\frac{1}{2}|\langle D^{2}(\tilde{u}^{n})_{\varepsilon}(\xi)(x-y),(x-y)\rangle|$
@@
On the other hand, calling $\delta_{2}=\delta_{2}(x)=\operatorname{dist}\big{(}x;\partial Q_{i}^{(1-\beta)\gamma}\big{)}$, taking the normalization $\nu(x)=\nabla(\tilde{u}^{n})_{\varepsilon}(x)/\|\nabla(\tilde{u}^{n})_{\varepsilon}(x)\|$, applying the change of variables $z=x-y$, exploiting the invariance of the integral on $B_{\delta_{2}}$ with respect to rotations, applying the change of variable $z\mapsto z\delta_{2}^{-1}$, and observing that, for every $\nu\in\mathbb{S}^{d-1}$,
$(1-s_{n})\int_{B_{1}}\frac{|z\cdot\nu|^{p}}{\|z\|^{d+s_{n}p}}\,\mathrm{d}z=(1-s_{n})\int_{0}^{1}\varrho^{p-s_{n}p-1}\,\int_{\partial B_{1}}|\eta\cdot\nu|^{p}\,\mathrm{d}\eta\,\mathrm{d}\varrho=K_{d,p}$ (3.23)
@@
[…]
@@
[…]
@@
$\begin{split}\lim_{\gamma\to 0^{+}}\sum_{i\in I_{\gamma}}K_{d,p}\int_{Q_{i}^{\gamma}}|\nabla\tilde{u}(x)|^{p}(f_{i}^{\gamma})^{2}\,\mathrm{d}x=K_{d,p}\|\nabla u\|^{p}_{p,f^{2}}\\ \lim_{\gamma\to 0^{+}}\sum_{i\in I_{\gamma}}K_{d,1}\int_{Q_{i}^{\gamma}}(f_{i}^{\gamma})^{2}\,\mathrm{d}|Du|=K_{d,1}\|Du\|_{1,f^{2}}.\end{split}$
@@
3.3. Proof of theorem 3.1(iii)
[…]
@@
$\begin{split}I_{1}&=\int_{\mathbb{R}^{d}}\int_{\{\|x-y\|\leq 1\}}\frac{|v(x)-v(y)|^{p}}{\|x-y\|^{d+sp}}f^{2}(y)\,\mathrm{d}x\,\mathrm{d}y,\\ I_{2}&=\int_{\mathbb{R}^{d}}\int_{\{\|x-y\|\leq 1\}}\frac{|v(x)-v(y)|^{p}}{\|x-y\|^{d+sp}}f(y)(f(x)-f(y))\,\mathrm{d}x\,\mathrm{d}y,\\ I_{3}&=\int_{\mathbb{R}^{d}}\int_{\{\|x-y\|\geq 1\}}\frac{|v(x)-v(y)|^{p}}{\|x-y\|^{d+sp}}f(x)f(y)\,\mathrm{d}x\,\mathrm{d}y.\end{split}$ (3.28)
@@
$\leq C(f,v)d\omega_{d}|(\operatorname{spt}{v})_{1}|\int_{0}^{1}\frac{1}{\varrho^{-p+sp}}\,\mathrm{d}\varrho,$
@@
$|v(x)-v(y)|^{p}$ $=|\nabla v(y)\cdot(x-y)|^{p}+O(\|x-y\|^{2p})$
$\leq|\nabla v(y)\cdot(x-y)|^{p}+\|D^{2}v\|_{\infty}^{2p}\,\|x-y\|^{2p}.$
@@
$\leq C(f,v)d\omega_{d}|(\operatorname{spt}{v})_{1}|\int_{0}^{1}\varrho^{p-1+p(1-s)}\,\mathrm{d}\varrho,$
@@
By the non-negativity of $I^{\prime\prime}_{1},I_{2},I_{3}$ and by (3.29)–(3.31), we thus get that
@@
Therefore, in order to conclude, we are left with showing that $\lim_{s\to 1^{-}}(1-s)I^{\prime}_{1}$ equals the right hand side of (3.26). Indeed, by the change of variables $z=x-y$, by setting $\nu(y)=\nabla v(y)/\|\nabla v(y)\|$, and using (3.23), we have
@@
[…]
Proof of theorem 3.1(iii).
Let $u\in W^{1,p}_{0}(\Omega)$ for $p>1$, or $u\in BV_{0}(\Omega)$ for $p=1$. By the density of $C_{c}^{\infty}(\Omega)$ in $W^{1,p}_{0}(\Omega)$ for $p>1$, or in $BV_{0}(\Omega)$ for $p=1$, we can find $(v^{k})_{k\in\mathbb{N}}\subset C_{c}^{\infty}(\Omega)$ such that $v^{k}\to u$ in $W^{1,p}_{0}(\Omega)$ for $p>1$, or in energy in $BV_{0}(\Omega)$ for $p=1$. In view of theorem 3.6, we thus get that
@@
3.4. Proof of theorem 1.1
We can now prove theorem 1.1 in full generality. Indeed, the result easily follows by combining theorem 3.1 with lemmas 3.7 and 3.8 below (under the same standing assumptions as stated at the very beginning of section 3).
@@
the conclusion immediately follows from (3.34) ∎
@@
$\begin{split}\int_{\mathbb{R}^{d}}\int_{B_{1}(y)^{c}}&\frac{|\tilde{u}^{n}(x)-\tilde{u}^{n}(y)|^{p}}{\|x-y\|^{d+s_{n}p}}\,\mathrm{d}x\,\mathrm{d}y\leq 2^{p}\int_{\mathbb{R}^{d}}\int_{B_{1}(y)^{c}}\frac{|\tilde{u}^{n}(x)|^{p}+|\tilde{u}^{n}(y)|^{p}}{\|x-y\|^{d+s_{n}p}}\,\mathrm{d}x\,\mathrm{d}y\\ &\leq 2^{p+1}\int_{\mathbb{R}^{d}}|\tilde{u}^{n}(y)|^{p}\int_{B_{1}(y)^{c}}\frac{\mathrm{d}x}{\|x-y\|^{d+s_{n}p}}\,\mathrm{d}y\\ &=2^{p+1}\|u^{n}\|^{p}_{p}\int_{1}^{+\infty}\frac{d\omega_{d}}{\varrho^{1+s_{n}p}}\,\mathrm{d}\varrho=\frac{C(d,p)\|u^{n}\|^{p}_{p}}{s_{n}}\leq C(d,p)\|u^{n}\|^{p}_{p}\end{split}$ (3.37)
@@
Since $f_{n}\to f$ uniformly and $f>0$, the constant $C$ can be made independent of $n$, and only on the uniform limit $f$. The conclusion hence follows by combining (3.37) and (3.38), and multiplying by $(1-s_{n})$, owing to the assumption (3.36). We omit the plain details. ∎
4. Proof of theorem 1.2

### v2

@@ 2.2. Compactness and characterization
$\|\tau_{h}\tilde{u}^{n}-\tilde{u}^{n}\|_{L^{p}(\mathbb{R}^{d})}\leq C(d,p,M,\Omega,f)\|h\|^{s_{n}},$ (3.14)
for all $h\in\mathbb{R}^{d}$ such that $\|h\|\leq 1$. Thus, owing to (3.1) and to (3.14), we can apply Theorem˜2.1 and find $u\in L^{p}(\mathbb{R}^{d})$ such that, up to a subsequence, $\tilde{u}^{n}\to u$ in $L^{p}(\mathbb{R}^{d})$. Furthermore, since $\tilde{u}^{n}=0$ for all $n\in\mathbb{N}$ on $\mathbb{R}^{d}\setminus\Omega$, we also have that $u=0$ on $\mathbb{R}^{d}\setminus\Omega$. Finally, letting $n\to\infty$ in (3.14), we have
$\|\tau_{h}u-u\|_{L^{p}(\mathbb{R}^{d})}\leq C(d,p,M,\Omega,f)\|h\|,$
for all $h$ with $\|h\|\leq 1$, so that $u\in W^{1,p}(\mathbb{R}^{d})$ for $p>1$ or $u\in BV(\mathbb{R}^{d})$ for $p=1$ by Theorem˜2.2. Since $u=0$ on $\mathbb{R}^{d}\setminus\Omega$ and $\Omega$ has Lipschitz boundary, we get that $u|_{\Omega}\in W^{1,p}_{0}(\Omega)$ for $p>1$, or $u|_{\Omega}\in BV(\Omega)$ for $p=1$, concluding the proof. ∎
3.2. Proof of Theorem˜3.1(ii)
@@
is finite. In addition, given $f\colon\mathbb{R}^{d}\to\mathbb{R}$, we let $f_{i}^{a}=\inf_{Q_{i}^{a}}f$. Notice that, whenever $f\in\mathrm{Lip}(\mathbb{R}^{d};(0,\infty))$, then $f_{i}^{a}>0$. Finally, we let $\eta\in C_{c}^{\infty}(B_{1})$ be such that $\eta\geq 0$ and $\int_{B_{1}}\eta\,\mathrm{d}x=1$ and, for every $\varepsilon>0$, we set $\eta_{\varepsilon}(\,\cdot\,)=\varepsilon^{-d}\eta(\,\cdot\,/\varepsilon)$. Accordingly, we let $u_{\varepsilon}=u\ast\eta_{\varepsilon}$ for every $\varepsilon>0$ and $u\in L^{1}_{\rm loc}(\mathbb{R}^{d})$.
@@
Let $p\in[1,\infty)$, $f\in\mathrm{Lip}_{b}(\mathbb{R}^{d};(0,\infty))$. There exist $\varepsilon,\beta,\gamma>0$ with $\varepsilon\ll\beta\ll\gamma$ such that
$\begin{split}\int_{Q_{i}^{(1-\beta)\gamma}}\int_{Q_{i}^{(1-\beta)\gamma}}&\frac{|\tilde{u}_{\varepsilon}(x)-\tilde{u}_{\varepsilon}(y)|^{p}}{\|x-y\|^{d+sp}}(f_{i}^{\gamma})^{2}\,\mathrm{d}x\,\mathrm{d}y\\ &\quad\qquad\leq\int_{Q_{i}^{\gamma}}\int_{Q_{i}^{\gamma}}\frac{|\tilde{u}(x)-\tilde{u}(y)|^{p}}{\|x-y\|^{d+sp}}f(x)f(y)\,\mathrm{d}x\,\mathrm{d}y\end{split}$ (3.15)
@@
for all $z\in B_{\varepsilon}$ and all $i\in\gamma\mathbb{Z}^{d}$. Indeed,
[…]
and therefore, to get (3.16), it is enough to choose $\varepsilon<\beta\gamma$ to ensure $(1-\beta)\gamma+\varepsilon<\gamma$. Using the definition of convolution, Jensen’s inequality, changing order of integration, performing the change of variables $x-z=\xi$ and $y-z=\zeta$, owing to (3.16), changing once again order of integration, using the fact that $\eta_{\varepsilon}$ has unit $L^{1}$ norm, and finally owing to the definition of $f_{i}^{\gamma}=\inf_{Q_{i}^{\gamma}}f$, we obtain
$\int_{Q_{i}^{(1-\beta)\gamma}}$ $\int_{Q_{i}^{(1-\beta)\gamma}}\frac{|\tilde{u}_{\varepsilon}(x)-\tilde{u}_{\varepsilon}(y)|^{p}}{\|x-y\|^{d+sp}}(f_{i}^{\gamma})^{2}\,\mathrm{d}x\,\mathrm{d}y$
$\leq\int_{Q_{i}^{(1-\beta)\gamma}}\int_{Q_{i}^{(1-\beta)\gamma}}\int_{B_{\varepsilon}}\frac{|\tilde{u}(x-z)-\tilde{u}(y-z)|^{p}}{\|x-z-(y-z)\|^{d+sp}}\eta_{\varepsilon}(z)(f_{i}^{\gamma})^{2}\,\mathrm{d}z\,\mathrm{d}x\,\mathrm{d}y$
$\leq\int_{B_{\varepsilon}}\int_{Q_{i}^{(1-\beta)\gamma}+z}\int_{Q_{i}^{(1-\beta)\gamma}+z}\frac{|\tilde{u}(\xi)-\tilde{u}(\zeta)|^{p}}{\|\xi-\zeta\|^{d+sp}}\eta_{\varepsilon}(z)(f_{i}^{\gamma})^{2}\,\mathrm{d}\xi\,\mathrm{d}\zeta\,\mathrm{d}z$
$\leq\int_{Q_{i}^{\gamma}}\int_{Q_{i}^{\gamma}}\frac{|\tilde{u}(\xi)-\tilde{u}(\zeta)|^{p}}{\|\xi-\zeta\|^{d+sp}}\,f(\xi)\,f(\zeta)\,\mathrm{d}\xi\,\mathrm{d}\zeta,$
@@
We are now ready to prove the $\liminf$ statement (ii) in Theorem˜3.1.
Proof of Theorem˜3.1(ii).
@@
otherwise inequality (3.2) is trivially true. Therefore, we can assume the validity of (3.1), which, in turn, implies that $u\in W^{1,p}_{0}(\Omega)$ for $p>1$, or $u\in BV(\Omega)$ for $p=1$. By Lemma˜3.5, there exist $\varepsilon,\beta,\gamma>0$ with $\varepsilon\ll\beta\ll\gamma$ such that
@@
$|\nabla(\tilde{u}^{n})_{\varepsilon}(x)\cdot(x-y)|$ $\leq|(\tilde{u}^{n})_{\varepsilon}(x)-(\tilde{u}^{n})_{\varepsilon}(y)|+\frac{1}{2}|\langle D^{2}(\tilde{u}^{n})_{\varepsilon}(\xi)(x-y),(x-y)\rangle|$
@@
On the other hand, calling $\delta_{2}=\delta_{2}(x)=\operatorname{dist}\big(x;\partial Q_{i}^{(1-\beta)\gamma}\big)$, taking the normalization $\nu(x)=\nabla(\tilde{u}^{n})_{\varepsilon}(x)/\|\nabla(\tilde{u}^{n})_{\varepsilon}(x)\|$, applying the change of variables z=xyz=x-y, exploiting the invariance of the integral on $B_{\delta_{2}}$ with respect to rotations, applying the change of variable $z\mapsto z\delta_{2}^{-1}$, and observing that, for every $\nu\in\mathbb{S}^{d-1}$
$(1-s_{n})\int_{B_{1}}\frac{|z\cdot\nu|^{p}}{\|z\|^{d+s_{n}p}}\,\mathrm{d}z=(1-s_{n})\int_{0}^{1}\varrho^{p-s_{n}p-1}\,\int_{\partial B_{1}}|\eta\cdot\nu|^{p}\,\mathrm{d}\eta\,\mathrm{d}\varrho=K_{d,p},$ (3.23)
@@
Multiplying the above inequality by $(1-s_{n})$, integrating over the cube $Q_{i}^{(1-\beta)\gamma}$ with respect to $x$, taking the $\liminf$ as $n\to\infty$, owing to Fatou’s lemma, we get that
@@
Thence, by taking the $\liminf$ as $n\to\infty$ of $(1-s_{n})$ times (3.17) and owing to (3.20)–(3.24), we have
@@
$\begin{split}\lim_{\gamma\to 0^{+}}\sum_{i\in I_{\gamma}}K_{d,p}\int_{Q_{i}^{\gamma}}|\nabla\tilde{u}(x)|^{p}(f_{i}^{\gamma})^{2}\,\mathrm{d}x=K_{d,p}\|\nabla u\|^{p}_{p,f^{2}}\\ \text{and}\quad\lim_{\gamma\to 0^{+}}\sum_{i\in I_{\gamma}}K_{d,1}\int_{Q_{i}^{\gamma}}(f_{i}^{\gamma})^{2}\,\mathrm{d}|Du|=K_{d,1}\|Du\|_{1,f^{2}}.\end{split}$
@@
3.3. Proof of Theorem˜3.1(iii)
We begin with the following preliminary result, establishing the $\limsup$ inequality (iii) in Theorem˜3.1 for smooth functions supported in $\Omega$.
@@
$\begin{split}I_{1}&=\int_{\mathbb{R}^{d}}\int_{\{\|x-y\|<1\}}\frac{|v(x)-v(y)|^{p}}{\|x-y\|^{d+sp}}f^{2}(y)\,\mathrm{d}x\,\mathrm{d}y,\\ I_{2}&=\int_{\mathbb{R}^{d}}\int_{\{\|x-y\|<1\}}\frac{|v(x)-v(y)|^{p}}{\|x-y\|^{d+sp}}f(y)(f(x)-f(y))\,\mathrm{d}x\,\mathrm{d}y,\\ I_{3}&=\int_{\mathbb{R}^{d}}\int_{\{\|x-y\|\geq 1\}}\frac{|v(x)-v(y)|^{p}}{\|x-y\|^{d+sp}}f(x)f(y)\,\mathrm{d}x\,\mathrm{d}y.\end{split}$ (3.28)
@@
$\leq C(f,v)d\omega_{d}|(\operatorname{spt}{v})_{1}|\int_{0}^{1}\frac{1}{\varrho^{-p+sp}}\,\mathrm{d}\varrho$
@@
$|v(x)-v(y)|^{p}\leq|\nabla v(y)\cdot(x-y)|^{p}+\|D^{2}v\|_{\infty}^{2p}\,\|x-y\|^{2p}.$
@@
$\leq C(f,v)d\omega_{d}|(\operatorname{spt}{v})_{1}|\int_{0}^{1}\varrho^{p-1+p(1-s)}\,\mathrm{d}\varrho$
@@
By the non-negativity of $I^{\prime\prime}_{1},I_{2},I_{3}$, and by (3.29)–(3.31), we thus get that
@@
Therefore, in order to conclude, we are left with showing that $\lim_{s\to 1^{-}}(1-s)I^{\prime}_{1}$ equals the right hand side of (3.26). Indeed, by the change of variables z=xyz=x-y, by setting $\nu(y)=\nabla v(y)/\|\nabla v(y)\|$, and using (3.23), we have
@@
Remark 3.7.
We underline that, in the chain of inequalities leading to (3.30), the Lipschitz regularity of the weight $f$ is crucial in order to ensure the finiteness of the integral in $\varrho$. If $f$ were only $\alpha$-Hölder, for some $\alpha\in(0,1)$, one would end up with $\varrho$ to the power $(\alpha-1-p+sp)$, which is not integrable in a neighborhood of the origin.
We are now ready to prove the $\limsup$ statement (iii) in Theorem˜3.1.
Proof of Theorem˜3.1(iii).
Let $u\in W^{1,p}_{0}(\Omega)$ for $p>1$, or $u\in BV(\Omega)$ for $p=1$. By the density of $C_{c}^{\infty}(\Omega)$ in $W^{1,p}_{0}(\Omega)$ for $p>1$, or in $BV(\Omega)$ for $p=1$, we can find $(v^{k})_{k\in\mathbb{N}}\subset C_{c}^{\infty}(\Omega)$ such that $v^{k}\to u$ in $W^{1,p}_{0}(\Omega)$ for $p>1$, or in energy in $BV(\Omega)$ for $p=1$. In view of Theorem˜3.6, we thus get that
@@
3.4. Proof of Theorem˜1.1
We can now prove Theorem˜1.1 in full generality. Indeed, the result easily follows by combining Theorem˜3.1 with Lemmas˜3.8 and 3.9 below (under the same standing assumptions as stated at the very beginning of Section˜3).
@@
the conclusion immediately follows from (3.34). ∎
@@
$\begin{split}\int_{\mathbb{R}^{d}}\int_{B_{1}(y)^{c}}&\frac{|\tilde{u}^{n}(x)-\tilde{u}^{n}(y)|^{p}}{\|x-y\|^{d+s_{n}p}}\,\mathrm{d}x\,\mathrm{d}y\leq 2^{p}\int_{\mathbb{R}^{d}}\int_{B_{1}(y)^{c}}\frac{|\tilde{u}^{n}(x)|^{p}+|\tilde{u}^{n}(y)|^{p}}{\|x-y\|^{d+s_{n}p}}\,\mathrm{d}x\,\mathrm{d}y\\ &\leq 2^{p+1}\int_{\mathbb{R}^{d}}|\tilde{u}^{n}(y)|^{p}\int_{B_{1}(y)^{c}}\frac{\mathrm{d}x}{\|x-y\|^{d+s_{n}p}}\,\mathrm{d}y\\ &=2^{p+1}\|u^{n}\|^{p}_{p}\int_{1}^{\infty}\frac{d\omega_{d}}{\varrho^{1+s_{n}p}}\,\mathrm{d}\varrho=\frac{C(d,p)\|u^{n}\|^{p}_{p}}{s_{n}}\leq C(d,p)\|u^{n}\|^{p}_{p}\end{split}$ (3.37)
@@
Since $f_{n}\to f$ uniformly and $f>0$, the constant $C$ can be made independent of $n$, and only dependent on the uniform limit $f$. The conclusion hence follows by combining (3.37) and (3.38), multiplying by $(1-s_{n})$, and owing to the assumption (3.36). We omit the plain details. ∎
4. Proof of Theorem˜1.2